# Algebraic dependence number and cardinality of generating iterated function systems

### v1

@@ 1. Introduction
which is called a self-similar set. We call $\{S_{i}\}_{i=1}^{m}$ a generating IFS for $K$. We say that the strong separation condition (SSC) holds for the IFS when the union is disjoint in (1.1), and such $K$ is called dust -like.
This paper is motivated by the inverse fractal problem in [11, 12, 20, 29]. Given a dust-like set, what can be said about its generating IFSs? This problem is also related to tiling theory and image compression, see [20, Section 1] for details. Our observation is that, the gap lengths of a dust-like set can determine its algebraic dependence number (Theorem 3.6), which gives a lower bound on the cardinality of the generating IFS (Corollary 3.8). In [15], the algebraic dependence number of an IFS is defined as the dimension over $\mathbb{Q}$ of the vector space generated by the logarithms of the contraction ratios of all similarities, minus 1. For a dust-like self-similar set in $\mathbb{R}^{n}$, [15, Theorem 5.7] shows that all its generating IFSs with the SSC have the same algebraic dependence number, so this number can be defined for a dust-like self-similar set as its intrinsic characteristic (independent of generating IFSs). This is done by considering certain measures on the dust-like self-similar set, while our characterisation is direct and intrinsic (without considering generating IFSs and measures), and we generalise this to compact metric spaces.
We also consider graph-directed IFSs and graph-directed attractors based on a given directed graph, which generalizes the concept of IFS and self-similar sets. It is introduced in [32], and have been studied a lot including the geometric aspects (see for example [8, 9, 14, 18, 22, 34, 38]) and fractal analysis [4, 6, 24, 33]. A directed graph $\left(V,E\right)$ consists of a finite set of vertices $V$ and a finite set of directed edges $E$ with loops and multiple edges allowed. Given a directed graph, we associate each vertex $v\in V$ with a metric space $(M^{v},d^{v})$. Let $E_{uv}\subset E$ be the set of edges from vertex $u$ to $v$. A graph-directed iterated function system (GD-IFS) $\left\{S_{e}:e\in E_{uv}\right\}$ consists of contracting similarities $S_{e}$ from $M_{v}$ to $M_{u}$, that is,
@@
see [13, Theorem 4.3.5 on p.128]. We call the above $\left(F_{u}\right)_{u\in V}$ the (list of) attractors of the GD-IFS, and each $F_{u}$ is called a GD-attractor. We say that the strong separation condition (SSC) holds for the GD-IFS, if the union is disjoint in (1.3) for each $u\in V$, and the GD-attractors are called dust-like.
The graph-directed setting appear frequently in dynamical systems and the study of self-similar sets. Certain complex dynamical systems can be regarded as conformal GD-IFSs using a Markov partition [16, Section 5.5]. The orthogonal projection of certain self-similar sets may be GD-attractors [17, Theorem 1.1]. When exact overlap occurs in an IFS, one can use a corresponding graph-directed system to study its attractor (see for example [10, 26]). Also in fractal analysis, GD-IFSs are used to determine whether the two walk-dimensions coincide for p.c.f. self-similar sets [23]. They are also related to tiling problems [36] and Automata [7].
Our results are useful for inhomogeneous (GD-)IFSs and (GD-)attractors, where the contraction ratios are not all equal. The idea is applying the ‘ratio analysis’ method in [28] to the ‘gap length set’ of a dust-like GD-attractor. Separation conditions are often required in fractal geometry to obtain some precise structure information, and relaxing them can be very difficult in many problems. In our paper, the SSC is (only) used for a precise formula of the gap length set [10, Theorem 2] for ratio analysis. We remark that the SSC is a common assumption for inhomogeneous IFSs in related problems (see for example [1, 15, 21]). In Corollary 3.10 we use the ‘full-measure’ condition as in [20, Theorem 4.1] to remove the SSC condition.
This paper is organised as follows. In Section 2, we recall the definition and expression of the gap length set of dust-like GD-attractors, and the required ratio analysis lemmas in previous studies. In Section 3, we present our results. In Section 3.1, we use ratio analysis on the gap length set to obtain some lemmas and the logarithmic commensurability for SSC generating IFSs (Theorem 3.2), which gives an alternative proof of [15, Theorem 5.7]. In Section 3.2, we present Theorem 3.6 which characterizes the algebraic dependence number of (GD-)IFS in terms of (the geometric sequences in) the gap length set. In Section 3.3, we present the lower bound estimate on the cardinality of generating IFS, with or without separation conditions (Corollaries 3.8 and 3.10).
@@ 2.1. Gap lengths
The concept and lemma in this subsection are from [10], which concerns the ‘gap sequence’ (see [35] for details in $\mathbb{R}^{n}$), a natural way to define the size and number of holes for disconnected sets. For a compact metric space $(K,d)$ and $\delta>0$, denote its diameter by diam$(K)$. We say that $x,y\in K$ are $\delta$-equivalent, if there is a $\delta$-chain connecting $x$ and $y$, that is, there are points $x_{i}\in K\ (i=1,2,\cdots,t)$ such that $x_{1}=x,x_{t}=y,d(x_{i},x_{i+1})\leq\delta$ for all $i$. According to this equivalence, we can divide $K$ into several $\delta$-equivalence classes (intuitively, these small pieces are separated by a distance $\delta$, while any two points in each piece can be connected by some $\delta$-chain). Let $\kappa(\delta)$ be the number of $\delta$-equivalence classes of $K$, which is finite due to the compactness of $K$ (since two points in an open $\delta/2$-metric ball are automatically $\delta$-equivalent, $\kappa(\delta)$ does not exceed the covering number of $K$ using open $\delta/2$-metric balls).
@@
For compact sets on $\mathbb{R}$, this definition is intuitive and coincides with [28, Definition 2.1], see explanation and an example (middle-third Cantor set) in [35, Section 1].
@@
The contraction ratio set of a GD-IFS is defined to be the set of the contraction ratios of the similarities, that is, $\{r_{e}:e\in E\}$ using the notion in (1.2).
The following Lemma is a direct corollary of [10, Theorem 2]. We define the product of $A,\ B\subset\mathbb{R}$ to be $AB=\{ab:\ a\in A,\ b\in B\}.$ We regard a real number as a set when encountering the product with a set in $\mathbb{R}$, and $AB$ is empty if $A$ is empty. Given a GD-IFS and a directed path $\mathbf{e}=e_{1}e_{2}\cdots e_{k}$ with edges $e_{i},$ (for which the terminal vertex of $e_{i}$ is the initial vertex of $e_{i+1}$ when $i=1,\mathbb{\cdots},k-1$), define $r_{\mathbf{e}}:=r_{e_{1}}r_{e_{2}}\cdots r_{e_{k}}$.
@@
When there is no directed path from $u$ to $v$, then $R_{uv}=\emptyset$. We mention that [10, Theorem 2] gives the construction of $\Lambda_{u},\ \Gamma_{u}$, although it is not used in our paper. Given a GD-IFS and its attractor list $(F_{u}\subset M^{u})_{u\in V}$, we obtain $\mathrm{GL}(F_{u})$ and define $\Gamma_{u}=\mathrm{GL}(F_{u})\bigcap[\delta,\infty)$ for all $u\in V$, where $\delta:=\inf\Big{\{}d^{u}(x_{e},y_{e}):\ x_{e}\in S_{e}(F_{v}),\ y_{e}\in F_{u}\setminus S_{e}(F_{v}),\ e\in E_{uv},\ u,\ v\in V\Big{\}}>0.$ Then define $\Lambda_{u}=\bigcup_{u\in V}\bigcup_{e\in E_{uv}}r_{e}(\Gamma_{v}\bigcap[\delta,r_{e}^{-1}\delta))$ for each $u\in V$, from which we know $\Lambda_{u}\subset(0,\delta)$.
We will use the following notions in [28, Definition 2.5] and a corollary of the above lemma.
@@ 2.2. Ratio analysis
In [28], the “ratio analysis” method is used to analyse sets $\Theta$ of positive real numbers, in terms of strictly decreasing geometric sequences $\{\theta^{\prime}r^{k}\}_{k=0}^{\infty}$ that are contained in $\Theta$. Whenever we mention ‘geometric sequence’ in this paper, we always assume it is infinitely-long.
@@
([28, Definition 2.4]) Let $\Theta\subset(0,\infty)$. For $\theta\in\Theta$, denote by
@@
This concept is quite natural since the gap length set of a dust-like attractor contain many geometric sequences. By definition, there is an obvious monotoniticy: when $\theta\in\Theta_{1}\subset\Theta_{2}\subset(0,\infty)$,
@@
([28, Lemma 2.6 (i)]) Let $A=\{a_{i}\}_{i=1}^{n}\subset(0,1)$ be a finite set, and $\lambda_{j}\ (j=1,\cdots,m)$ be positive real numbers (not necessarily distinct). Let $\Theta=\mathop{\textstyle\bigcup}\limits_{j=1}^{m}\lambda_{j}A_{j}$ where $A_{j}\subset A^{\mathbb{Z}_{+}}$ for $1\leq j\leq m$. Then $R_{\Theta}(\theta)\subset A^{\mathbb{Q}_{+}^{\ast}}$ for all $\theta\in\Theta$.
@@
([28, Corollary 2.7]) Let $X\subset(0,1)$ and $\Lambda\subset(0,\infty)$ be two finite sets. Then
@@ 3.1. Using ratio analysis to obtain logarithmic commensurability
[19, Conjecture 1.2] Suppose that $K,\ F$ are two totally disconnected non-singleton self-similar sets in $\mathbb{R}^{n}$, which are the attractors of two IFS with contraction ratio sets $X,\ Y$ respectively. Suppose that there exists an affine map $f$ in $\mathbb{R}^{n}$ such that $f(F)\subset K,$ then the logarithmic commensurability holds: $Y\subset X^{\mathbb{Q}_{+}^{*}}.$
@@ 3.2. Intrinsic characterisation of algebraic dependence number
Given a GD-IFS $\digamma$ with contraction ratio set $A$, define its algebraic independence number as the dimension of the vector space $\mathrm{span}\log A$, and its algebraic dependence number as its algebraic independence number minus 1.
@@
Suppose that $\digamma$ is a GD-IFS satisfying the SSC with attractor list $(F_{u}\subset M^{u})_{u\in V}$ (on complete metric spaces), based on a directed graph $(V,E)$ with $d_{w}\geq 2$ for all $w\in V$, then its algebraic independence number is no less than the dimension of
@@
In fact we prove a stronger property (3.5) for dust-like self-similar sets: the vector space generated by the logarithms of the contraction ratios of any SSC generating IFS (over $\mathbb{Q}$) is equal to the vector space generated by the logarithms of all the common ratios of geometric sequences in the gap length set (over $\mathbb{Q}$).
@@ 3.3. Lower bound on the cardinality of generating IFS
In particular, for a dust-like self-similar set $K\subset M$ (where $M$ is a complete metric space), the cardinality of each generating IFS satisfying the SSC is no less than dimension of $\mathrm{span}\Big{\{}\mathrm{log}\,r:\ r\in\bigcup_{\theta\in\mathrm{GL}(K)}R_{\mathrm{GL}(K)}(\theta)\Big{\}}=\mathrm{span}\{\mathrm{log}\,r:r\in R_{\mathrm{GL}(K)}(\theta)\}$ for any small enough $\theta\in\mathrm{GL}(K)$.
@@
Furthermore, if the dust-like self-similar set is on $\mathbb{R}$ and is full-measure, we can remove the separation condition (SSC) on the generating IFS. We say that a nonempty compact set $K$ is full-measure if $\mathcal{H}^{\mathrm{dim_{H}}K}(K)=(\mathrm{diam}(K))^{\mathrm{dim_{H}}K}.$ ($\mathrm{dim_{H}}$ is the Hausdorff dimension)
@@
Suppose that a full-measure set $K\subset\mathbb{R}$ is the attractor of an IFS $\Phi$ satisfying the SSC, then the cardinality of each generating IFS of $K$ is no less than dimension of $\mathrm{span}\Big{\{}\mathrm{log}\,r:\ r\in\bigcup_{\theta\in\mathrm{GL}(K)}R_{\mathrm{GL}(K)}(\theta)\Big{\}}$. If further, the logarithm of the contraction ratio of each contracting similarity in $\Phi$ is different from each other and linearly independent over $\mathbb{Q}$, then $\Phi$ has the minimal cardinality among all generating IFSs of $K$.
@@
For any generating IFS $\{S_{i}\}_{i=1}^{m}$ of $K$, there exists $I\in\{1,2,\cdots,m\}$ such that, the sub-IFS $\Psi:=\{S_{i}\}_{i\in I}$ also has attractor $K$ and satisfies the SSC. Indeed, whenever $S_{i}(K)\bigcap S_{j}(K)\neq\emptyset$, by [20, The claim in the proof of Theorem 4.1], $S_{i}(K)\subset S_{j}(K)$ (or $S_{i}(K)\supset S_{j}(K)$), so one can remove $i$ (or $j$) from $\{1,2,\cdots,m\}$, and repeat this process until the SSC is satisfied. The first inclusion then follows by using Corollary 3.8 for the lower bound of #$\#\Psi$ (also a lower bound for $m$).
For the second inclusion, by assumption, the algebraic independence number of $\Phi$ is exactly its cardinality #$\#\Phi$, thus the algebraic independence number of $\Psi$ is also #$\#\Phi$ by [15, Theorem 5.7]. Since the cardinality of $\Psi$ is no less than the algebraic independence number of $\Psi$, which is #$\#\Phi$, we obtain the desired by noting $m\geq\#\Psi$. $\square$
The full-measure condition is easy to verify by the similarities of an IFS on $\mathbb{R}$, see [20, Remark 4.1] for a practical sufficient and necessary condition and examples. We end our paper by a brief overview on the full-measure condition. It is firstly discussed in Hausdorff’s paper [25], then extended by Marion (called ‘perfect isotopic’) [30, 31], and independently by Ayer and Strichartz on self-similar sets (several years later): they give the full-measure criterion for dust-like self-similar sets on $\mathbb{R}$. There is no known analogous criterion in higher dimensional case to the best of our knowledge. The study of full-measure condition for a GD-attractor is more complicated than that for self-similar set (see [5, Theorem 4.6]), and also, there is no known analogous criterion even on $\mathbb{R}$.
Acknowledgement. The author was supported by National Natural Science Foundation of China (12271282).

### v3

@@ 1. Introduction
which is called a self-similar set. We call $\{S_{i}\}_{i=1}^{m}$ a generating IFS of $K$. We say that the strong separation condition (SSC) holds for the IFS when the union is disjoint in (1.1), and such $K$ is called dust -like.
This paper is motivated by the inverse fractal problem in [11, 12, 21, 29]. Given a dust-like set, what can be said about its generating IFSs? This problem is also related to tiling theory and image compression, see [21, Section 1] for details. Our results reveal the connection between the gap lengths of a dust-like set and its algebraic dependence number (Theorem 3.6), which can be used to give a lower bound on the cardinality of the generating IFS (Corollary 3.8). The definition of gap lengths is intrinsic, and it is simply the lengths of the disjoint (non-empty bounded) complementary intervals for a dust-like set on $\mathbb{R}$.
In [15], the algebraic dependence number of an IFS is defined as the dimension over $\mathbb{Q}$ of the vector space generated by the logarithms of the contraction ratios of all similarities, minus 1. For a dust-like self-similar set in $\mathbb{R}^{n}$, [15, Theorem 5.7] shows that all its generating IFSs with the SSC have the same algebraic dependence number, so this number can be defined for a dust-like self-similar set as its intrinsic characteristic (independent of generating IFSs). This is done by considering certain measures on the dust-like self-similar set, while our characterisation is direct and intrinsic (without considering generating IFSs and measures), and we generalise this to compact metric spaces. To be more detailed, our Theorem 3.6 gives an intrinsic quantitative characterisation of this number: it is the dimension over $\mathbb{Q}$ of the vector space generated by the logarithms of all the common ratios of infinite geometric sequences in the gap length set (the collection of all gap lengths), minus 1.
We also consider graph-directed IFSs and graph-directed attractors based on a given directed graph, which generalizes the concept of IFS and self-similar sets. It is introduced in [32] and has been studied intensively, including the geometric aspects (see for example [8, 9, 14, 19, 23, 34, 38]) and fractal analysis [4, 6, 25, 33]. A directed graph $\left(V,E\right)$ consists of a finite set of vertices $V$ and a finite set of directed edges $E$ with loops and multiple edges allowed. Given a directed graph, we associate each vertex $v\in V$ with a metric space $(M^{v},d^{v})$. Let $E_{uv}\subset E$ be the set of edges from vertex $u$ to $v$. A graph-directed iterated function system (GD-IFS) $\left\{S_{e}:e\in E_{uv}\right\}$ consists of contracting similarities $S_{e}$ from $M_{v}$ to $M_{u}$, that is,
@@
see [13, Theorem 4.3.5 on p.128]. We call the above $\left(F_{u}\right)_{u\in V}$ the (list of) attractors of the GD-IFS, and each $F_{u}$ is called a GD-attractor. We say that the strong separation condition (SSC) holds for the GD-IFS, if the union is disjoint in (1.3) for each $u\in V$, and the GD-attractors are called dust-like. Clearly, an IFS and its attractor can be viewed as a GD-IFS and its GD-attractor when #$\#V=1$, where ‘##\##’ denotes the cardinality of a finite set throughout the paper.
The graph-directed setting appear frequently in dynamical systems and the study of self-similar sets. Certain complex dynamical systems can be regarded as conformal GD-IFSs using a Markov partition [16, Section 5.5]. The orthogonal projection of certain self-similar sets may be GD-attractors [18, Theorem 1.1]. When exact overlap occurs in an IFS, one can use a corresponding graph-directed system to study its attractor (see for example [10, 27]). Also in fractal analysis, GD-IFSs are used to determine whether the two walk-dimensions coincide for p.c.f. self-similar sets [24]. They are also related to tiling problems [36] and Automata [7].
The idea is applying the ratio analysis method in [17] to the gap length set of a dust-like GD-attractor, which exhibits a new application of the ratio analysis method and the gap sequence (see [35] for applications to the Lipschitz equivalence problem and box dimension estimate). Our results are useful for inhomogeneous (GD-)IFSs and (GD-)attractors, where the contraction ratios are not all equal. Separation conditions are often required in fractal geometry to obtain some precise structure information, and relaxing them can be very difficult in many problems. In our paper, the SSC is (only) used for a precise formula of the gap length set [10, Theorem 2] for ratio analysis. We remark that the SSC is a common assumption for inhomogeneous IFSs in related problems (see for example [1, 15, 22]). In Corollary 3.10 we use the ‘full-measure’ condition as in [21, Theorem 4.1] to remove the SSC condition.
This paper is organised as follows. In Section 2, we introduce the definition and expression of the gap length set of a dust-like GD-attractor, and the required ratio analysis lemmas in [17]. In Section 3, we present our results. In Section 3.1, we use ratio analysis on the gap length set to obtain some lemmas and the logarithmic commensurability for SSC generating IFSs (Theorem 3.2), which gives an alternative proof of [15, Theorem 5.7]. In Section 3.2, we present Theorem 3.6, which characterizes the algebraic dependence number of (GD-)IFS in terms of (the geometric sequences in) the gap length set. In Section 3.3, we present the lower bound estimate on the cardinality of generating IFS, with or without separation conditions (Corollaries 3.8 and 3.10).
@@ 2.1. Gap lengths
The concept and lemma in this subsection are from [10], which concerns the gap sequence, a natural way to define the size and number of holes for disconnected sets (see [35] for details on $\mathbb{R}^{n}$). For a compact metric space $(K,d)$ and $\delta>0$, denote its diameter by diam$(K)$. We say that $x,y\in K$ are $\delta$-equivalent, if there is a $\delta$-chain connecting $x$ and $y$, that is, there are points $x_{i}\in K\ (i=1,2,\cdots,t)$ such that $x_{1}=x,x_{t}=y,d(x_{i},x_{i+1})\leq\delta$ for all $i$. According to this equivalence, we can divide $K$ into several $\delta$-equivalence classes (intuitively, these small pieces are separated by a distance $\delta$, while any two points in each piece can be connected by some $\delta$-chain). Let $\kappa(\delta)$ be the number of $\delta$-equivalence classes of $K$, which is finite due to the compactness of $K$ (since two points in an open $\delta/2$-metric ball are automatically $\delta$-equivalent, $\kappa(\delta)$ does not exceed the covering number of $K$ using open $\delta/2$-metric balls).
@@
For compact sets on $\mathbb{R}$, this definition is intuitive and coincides with [17, Definition 2.1], see explanation and an example (middle-third Cantor set) in [35, Section 1].
@@
The contraction ratio set of a GD-IFS is defined to be the set of the contraction ratios of the similarities, that is, $\{r_{e}:e\in E\}$ using the notation in (1.2).
The following Lemma is a direct corollary of [10, Theorem 2]. We define the product of $A,\ B\subset\mathbb{R}$ to be $AB=\{ab:\ a\in A,\ b\in B\}.$ We regard a real number as a set when encountering the product with a set in $\mathbb{R}$, and $AB$ is empty if $A$ is empty. Given a GD-IFS and a directed path $\mathbf{e}=e_{1}e_{2}\cdots e_{k}$ with edges $e_{i}$ (for which the terminal vertex of $e_{i}$ is the initial vertex of $e_{i+1}$ when $i=1,\mathbb{\cdots},k-1$), define $r_{\mathbf{e}}:=r_{e_{1}}r_{e_{2}}\cdots r_{e_{k}}$.
@@
When there is no directed path from $u$ to $v$, then $R_{uv}=\emptyset$. We mention that [10, Theorem 2] gives the construction of $\Lambda_{u}$ and $\Gamma_{u}$, although it is not used in our paper. Given a GD-IFS and its attractor list $(F_{u}\subset M^{u})_{u\in V}$, we obtain $\mathrm{GL}(F_{u})$ and define $\Gamma_{u}=\mathrm{GL}(F_{u})\bigcap[\delta,\infty)$ for all $u\in V$, where $\delta:=\inf\Big{\{}d^{u}(x_{e},y_{e}):\ x_{e}\in S_{e}(F_{v}),\ y_{e}\in F_{u}\setminus S_{e}(F_{v}),\ e\in E_{uv},\ u,\ v\in V\Big{\}}>0$ due to the SSC. Then define $\Lambda_{u}=\bigcup_{u\in V}\bigcup_{e\in E_{uv}}r_{e}(\Gamma_{v}\bigcap[\delta,r_{e}^{-1}\delta))$ for each $u\in V$, from which we know $\Lambda_{u}\subset(0,\delta)$.
We will use the following notions in [17, Definition 2.5] and a corollary of the above lemma.
@@ 2.2. Ratio analysis
In [17], the ratio analysis method is used to analyse sets $\Theta$ of positive real numbers, in terms of strictly decreasing geometric sequences $\{\theta^{\prime}r^{k}\}_{k=0}^{\infty}$ that are contained in $\Theta$. Whenever we mention ‘geometric sequence’ in this paper, we always assume it is infinitely-long.
@@
([17, Definition 2.4]) Let $\Theta\subset(0,\infty)$. For $\theta\in\Theta$, denote by
@@
This concept is quite natural since the gap length set of a dust-like attractor contains many geometric sequences. By definition, there is an obvious monotoniticy: when $\theta\in\Theta_{1}\subset\Theta_{2}\subset(0,\infty)$,
@@
([17, Lemma 2.6 (i)]) Let $A=\{a_{i}\}_{i=1}^{n}\subset(0,1)$ be a finite set, and $\lambda_{j}\ (j=1,\cdots,m)$ be positive real numbers (not necessarily distinct). Let $\Theta=\mathop{\textstyle\bigcup}\limits_{j=1}^{m}\lambda_{j}A_{j}$ where $A_{j}\subset A^{\mathbb{Z}_{+}}$ for $1\leq j\leq m$. Then $R_{\Theta}(\theta)\subset A^{\mathbb{Q}_{+}^{\ast}}$ for all $\theta\in\Theta$.
@@
([17, Corollary 2.7]) Let $X\subset(0,1)$ and $\Lambda\subset(0,\infty)$ be two finite sets. Then
@@ 3.1. Using ratio analysis to obtain logarithmic commensurability
[20, Conjecture 1.2] Suppose that $K,\ F$ are two totally disconnected non-singleton self-similar sets in $\mathbb{R}^{n}$, which are the attractors of two IFSs with contraction ratio sets $X,\ Y$ respectively. Suppose that there exists an affine map $f$ in $\mathbb{R}^{n}$ such that $f(F)\subset K,$ then the logarithmic commensurability holds: $Y\subset X^{\mathbb{Q}_{+}^{*}}.$
@@ 3.2. Intrinsic characterisation of algebraic dependence number
Given a GD-IFS with contraction ratio set $A$, define its algebraic independence number as the dimension of the vector space $\mathrm{span}\log A$, and its algebraic dependence number as its algebraic independence number minus 1.
@@
Suppose that $\digamma$ is a GD-IFS satisfying the SSC with GD-attractors $(F_{u}\subset M^{u})_{u\in V}$ (on complete metric spaces), based on a directed graph $(V,E)$ with $d_{w}\geq 2$ for all $w\in V$, then its algebraic independence number is no less than the dimension of
@@
In fact we prove a stronger property (3.5) for dust-like self-similar sets: the vector space generated by the logarithms of the contraction ratios of any SSC generating IFS (over $\mathbb{Q}$) is exactly the vector space generated by the logarithms of all the common ratios of geometric sequences in the gap length set (over $\mathbb{Q}$).
@@ 3.3. Lower bound on the cardinality of generating IFS
In particular, for a dust-like self-similar set $K\subset M$ (where $M$ is a complete metric space), the cardinality of its generating IFS satisfying the SSC is no less than dimension of
$\mathrm{span}\Big{\{}\mathrm{log}\,r:\ r\in\bigcup_{\theta\in\mathrm{GL}(K)}R_{\mathrm{GL}(K)}(\theta)\Big{\}}=\mathrm{span}\{\mathrm{log}\,r:r\in R_{\mathrm{GL}(K)}(\theta)\}$
for any small enough $\theta\in\mathrm{GL}(K)$.
@@
Furthermore, if the dust-like self-similar set is on $\mathbb{R}$ and is full-measure, we can remove the separation condition (SSC) on the generating IFS. We say that a nonempty compact set $K$ is full-measure if $\mathcal{H}^{\mathrm{dim_{H}}K}(K)=(\mathrm{diam}(K))^{\mathrm{dim_{H}}K}$ (where $\mathrm{dim_{H}}$ is the Hausdorff dimension).
@@
Suppose that a full-measure set $K\subset\mathbb{R}$ is the attractor of an IFS $\Phi$ satisfying the SSC, then the cardinality of every generating IFS of $K$ is no less than the dimension of $\mathrm{span}\Big{\{}\mathrm{log}\,r:\ r\in\bigcup_{\theta\in\mathrm{GL}(K)}R_{\mathrm{GL}(K)}(\theta)\Big{\}}$. If further, the logarithm of the contraction ratio of each contracting similarity in $\Phi$ is different from each other and linearly independent over $\mathbb{Q}$, then $\Phi$ has the minimal cardinality among all generating IFSs of $K$.
@@
For any generating IFS $\{S_{i}\}_{i=1}^{m}$ of $K$, there exists $I\in\{1,2,\cdots,m\}$ such that, the sub-IFS $\Psi:=\{S_{i}\}_{i\in I}$ also has attractor $K$ and satisfies the SSC, and
$m\geq\#\Psi.$ (3.6)
Indeed, whenever $S_{i}(K)\bigcap S_{j}(K)\neq\emptyset$, by [21, The claim in the proof of Theorem 4.1], $S_{i}(K)\subset S_{j}(K)$ (or $S_{i}(K)\supset S_{j}(K)$), so one can remove $i$ (or $j$) from $\{1,2,\cdots,m\}$, and repeat this process until the SSC is satisfied. The first inclusion then follows by using Corollary 3.8 for the lower bound of #$\#\Psi$, which is also a lower bound for $m$ by (3.6).
For the second inclusion, by assumption, the algebraic independence number of $\Phi$ is exactly its cardinality #$\#\Phi$, thus the algebraic independence number of $\Psi$ is also #$\#\Phi$ by [15, Theorem 5.7]. Since the cardinality of $\Psi$ is no less than the algebraic independence number of $\Psi$, which is #$\#\Phi$, we obtain $m\geq\#\Phi$ by (3.6), showing the desired. $\square$
The full-measure condition is easy to verify by the similarities of an IFS on $\mathbb{R}$, see [21, Remark 4.1] for a practical sufficient and necessary condition and examples. We end our paper by a brief overview on the full-measure condition. It is firstly discussed in Hausdorff’s paper [26], then extended by Marion (called ‘perfect isotopic’) [30, 31], and independently by Ayer and Strichartz [3] on self-similar sets (several years later): they give the full-measure criterion for dust-like self-similar sets on $\mathbb{R}$. There is no known analogous criterion in higher dimensional case to the best of our knowledge. The study of full-measure condition for a GD-attractor is more complicated than that for self-similar set (see [5, Theorem 4.6]), and there is no known analogous criterion even on $\mathbb{R}$ to the best of our knowledge.
Acknowledgement. The author thanks Kenneth Falconer for valuable suggestions.